What is the x intercept for y= -3/4x+5. I have (6.75,0)which is wrong?
Found 2 solutions by Alan3354, edjones:
Answer by Alan3354(69443)

What is the x intercept for y= -3/4x+5. I have (6.75,0)which is wrong?
-------------
At the x-intercept, y = 0
--------
0 = -(3/4)x + 5
3x/4 = 5
3x = 20
x = 20/3

Answer by edjones(8007)

y= -3/4x+5
(-3/4)x+5=0
(-3/4)x=-5
x=-5 / -3/4
x=-5 * -4/3
=20/3 you weren't wrong by much.
.
Ed